OpenStudy (anonymous):

2/5 , 3/5 which rational number in each pair is greater

OpenStudy (anonymous):

3/5

OpenStudy (anonymous):

3/5=.6 2/5=.4

OpenStudy (anonymous):

3/5

OpenStudy (amistre64):

i have to wonder where the "study" part comes into play with these types of replies ....

jhonyy9 (jhonyy9):

yes amistre right very

jhonyy9 (jhonyy9):

tuttordiffeq how you have wrote 3/5=6 so than 3=6*5=30 this is correct ? - or 2/5=4 so than 2=4*5=20 this is right ?

jhonyy9 (jhonyy9):

so 3/5=0,6 and 2/5=0,4 what is greater ?

OpenStudy (amistre64):

 which is more; 2 out of 5 pieces of the pie? or 3 out of 5 pieces of the pie?